# Quantum mechanics provides the physical basis of teleological evolutions

### v3

@@ 1 Introduction
Of course, while the teleological character of quantum algorithms, as we shall see, should be certain, the fact that—under the quantum cosmological assumption—it would extend to the evolutions of the living should be considered, at most, as plausible as the quantum cosmological assumption. Anyway, all this should be quite different from the present widespread belief that the notion of a teleological evolution of the living is an absurdity. At least, it should cast a doubt on the present position of science.
Let us eventually note that, so far, holism and finalism have been the enemy of the reductionist and mechanistic vision of science. In the present theory of the quantum computational speedup, they would acquire a physical basis. Of course, this would not touch on the immense practical value of the mechanistic and reductionist view. That the description of the evolution of any system is the trace over all the rest of the quantum description of the evolution of the universe, obviously, holds only in principle: in practice, we cannot have such descriptions. However, this contrast between the quantum and the classical view might pinpoint a direction of the research. At a toy model, or theoretical, level, it could be interesting to study how this “tracing over all the rest” conciliates a teleological character of the evolution of a system in the longer term with a classical character in the shorter term. This might even be a further perspective under which to study the unification between quantum and classical mechanics.
@@ 2.7 About the teleological character of the evolutions of the living
We have already seen in the Introduction that, under the quantum cosmological assumption and for the Fine-tuned Universe version of the Anthropic Principle, the physical basis of the teleological character of quantum algorithms could also apply to the evolution of the universe toward the value of life. This section expands on some facts that, for compactness, were just mentioned in the Introduction.
Of course, in the universe we are in, the possible teleological character of the evolution of the universe toward the value becomes the possible teleological character of the evolution of the universe toward life. Then, it is easily seen that this possible teleological character must be inherited by the evolution of life on Earth. In principle, the latter evolution can be derived from the evolution of the universe toward life by tracing all the rest. Conversely, we can speak of an image of the evolution of life on Earth in the evolution of the universe toward life. Then, first, evolution times must be the same for the evolution of life on Earth and its image in the evolution of the universe. Second, the amount of physical resources used by the evolution of life on Earth cannot exceed the amount of physical resources used by its image in the evolution of the universe. In fact, the latter resources comprise the former.
@@
Not to hide behind a finger, let us exemplify what a teleological character of the evolution of life would imply; it will be very destabilizing with respect to the present way of viewing the evolutions of the living.
Let us begin with the evolution of species. It is known that, for many millions of years, the ancestors of the birds were feathered dinosaurs who did not fly. Very light, strong but flexible, cave and air-proof, even at that time feathers were ideal for flight, very close to those of the future birds. So why anticipate something designed for flight many millions of years in advance?
@@
All this could remain true in the teleological view with just one, but very significant, addition: Those random mutations that would have led in the future to the conquest of the air would have more likely occurred. See our preprint [6] for a more comprehensive discussion of the teleological notion.
All this is obviously destabilizing; it is a way of viewing the time-development of macroscopic reality that goes against the current scientific (causal) way of viewing it. Indeed, it revamps the natural philosophy views that were ascribed by science to the category of the unscientific beliefs.
To the above example of teleological evolution of life, which concerned the evolution of the DNA, we should add another example that concerns the environment of the evolution of life. For example, what we humans are today has been affected by the drying up of an area of Africa in which our ancestors lived on trees three million years ago. Of course, there is nothing destabilizing in this. However, the time-symmetric fact that what we humans are today is also affected by what will concern our descendants in three million years’ time might be very difficult to digest. Indeed, it would profoundly change our (causal) way of viewing how things evolve in time.
Of course, all this is under a quantum cosmological reading of the Fine-tuned Universe version of the Anthropic Principle and is thus hypothetical. What we would consider certain is the teleological character of quantum algorithms.
Things could be further explored by investigating, either theoretically or by means of toy models, the character of the “tracing over all the rest”. We should consider the unitary quantum evolution of a system endowed with quantum speedup, standing for the evolution of the universe toward life, and derive from it the evolution of a subsystem of it, standing for the evolution of life on Earth, by tracing over all the rest. We should look for a way of obtaining an evolution of the subsystem that is teleological in the longer term and (about) classical in the shorter term. If this were possible, of course, it would go in favor of the above quantum cosmological reading of that version of the Anthropic Principle. By the way, it might also be a way of addressing the problem of unifying quantum and classical theory.
Another question is how to describe a macroscopic evolution endowed with quantum speedup. Of course, classical physics would turn out to be unable to describe not only the processes of the microscopic world but also the teleological evolutions of the macroscopic world.
@@
Let us end the present section by recalling what, to our knowledge, are the key steps that led to the discovery of the quantum computational speedup. According to the present theory, this discovery would be even more fundamental than currently thought. The first step was showing that computation is possible in the quantum world and introducing the fundamental notion of quantum bit, by Finkelstein [19] in 1969; that was also the first time that quantum physics moved to the world of the complex, indeed to a cellular automaton model of the universe—this was in the context of an attempt to unify relativity and quantum theory. Then, there was the introduction of the notion of logically reversible computation by Bennett [20] in 1973, of course, a prerequisite for the building up of computations in unitary quantum evolutions. The discovery of the superior computational efficiency of the quantum world with respect to that of the classical world in 1982 is due to Feynman [21]. Still in 1982, there has been the introduction of the notion of conservative logic by Fredkin and Toffoli [22]: with it, we have the logically reversible gates needed to perform quantum computations. The seminal quantum algorithm that delivers a quantum speedup was conceived by Deutsch [23] in 1985.
@@ 4 Conclusions
The flip side of the destabilizing character of the work is its innovative character. Let us distinguish between what should be certain and what should be plausible. About what should be the certain: (i) The work provides a unified quantitative explanation of the quantum speedups that is still missing from mainstream quantum mechanics. (ii) It provides a theoretical confirmation of the incompleteness of the quantum description of quantum nonlocality sustained by Einstein and others, also providing the local hidden variables envisaged by those authors to complete it. (iii) It highlights the teleological character of quantum algorithms; given the negative view that the teleological notion has in today’s science and the scope of this notion, this should be the most important result. About what is just plausible: It shows that, under a quantum cosmological reading of the Fine-tuned Universe version of the Anthropic Principle, the physical basis of the teleological character of quantum algorithms applies as well to the evolutions of the living.

### v4

@@ 1 Introduction
Note that the fact that a teleological evolution is also an evolution endowed with a quantum speedup is a novelty for the teleological notion. It could contribute to explaining steps of natural evolutions that are too rapid to be explained by the progressive character of Darwinian evolutions.
Summing up, while the teleological character of quantum algorithms, as we shall see, should be certain, the fact that—under the quantum cosmological assumption—it would extend to the evolutions of the living should be considered, at most, as plausible as the quantum cosmological assumption. Anyway, all this should be quite different from the present widespread belief that the notion of a teleological evolution of the living is an absurdity. At least, it should cast a doubt on this position of science.
Let us eventually note that, so far, holism and finalism have been the enemy of the reductionist and mechanistic vision of science. In the present theory of the quantum computational speedup, they would acquire a physical basis. Of course, this would not touch on the immense practical value of the mechanistic and reductionist view. That the description of the evolution of any system is the trace over all the rest of the quantum description of the evolution of the universe holds only in principle: obviously, in practice, we cannot have such descriptions. However, this contrast between the quantum and the classical view might pinpoint a direction of the research. At a toy model, or theoretical, level, it could be interesting to study how this “tracing over all the rest” conciliates a teleological character of the evolution of a system (standing for the evolution of life on Earth) in the longer term with a classical character in the shorter term. This might even be a further perspective under which to study the possibility of unifying quantum and classical mechanics.
@@ 2.7 About the teleological character of the evolutions of the living
We have already seen in the Introduction that, under the quantum cosmological assumption and for the Fine-tuned Universe version of the Anthropic Principle, the physical basis of the teleological character of quantum algorithms could also apply to the evolution of the universe toward the value of life. This section expands on some facts that, for compactness, were just mentioned, or omitted, in the Introduction.
Let us show that the possible teleological character of the evolution of the universe toward life must be inherited by the evolution of life on Earth. In principle, the latter evolution can be derived from that of the universe toward life by tracing out all the rest. Conversely, we can speak of an image of the evolution of life on Earth in the evolution of the universe toward life. Then, first, evolution times must be the same for the evolution of life on Earth and its image in the evolution of the universe. Second, the amount of physical resources used by the evolution of life on Earth cannot exceed the amount of physical resources used by its image in the evolution of the universe. In fact, the latter resources comprise the former.
@@
Not to hide behind a finger, let us exemplify what a teleological character of the evolution of life would imply; of course, it will be very destabilizing with respect to the present way of viewing both the evolutions of the living and physical reality.
Let us begin with the evolution of species. It is known that, for many millions of years, the ancestors of the birds were feathered dinosaurs who did not fly. Very light, strong but flexible, cave and air-proof, from the fossils found, even at that time feathers were ideal for flight, very close to those of the future birds. So, why anticipate something designed for flight many millions of years in advance?
@@
All this could remain true in the teleological view with just one, but very significant, addition: Those random mutations that would have led in the future to the conquest of the air would have more likely occurred. In our possibly biased opinion, this mutual causality explanation of the anticipation in question would be much more convincing than the customary causal one. See our preprint [6] for a more comprehensive discussion of the teleological notion.
Another possible consequence is for the fact that, according to the present theory, a teleological evolution is an evolution endowed with quantum speedup. This could contribute to explaining why some steps of the evolutions of the living have been so rapid to defy Darwinian gradualism, see [18]. Have been of this kind, in particular, the Cambrian Explosion and modern ecological adaptation.
Let us also recall what, here, is a very interesting observation of the French biologist Rémy Chauvin [19]. He pointed out the possibility that some strangely fast evolutionary changes of ants could happen before their being evolutionary advantages.
As is well known, this biologist merged important contributions to mainstream science with unconventional positions much criticized by colleagues. In addition to highlighting the possibility that some evolutionary changes are teleological in character, another unconventional position was his public interest in parapsychology (he wrote books on the subject). Note the parallel with the present situation. Let us consider the advance knowledge, on the part of the problem-solver, of any of the possible halves of the information about the solution of the problem she will read in the future. Besides being a completely technical notion at the basis of the present teleological notion (the problem-solver is a logical agent), under the quantum cosmological assumption it could well be the physical basis of the parapsychological notion of precognition.
Of course, being under the quantum cosmological assumption, all the above is hypothetical. This is to decouple it from the teleological character of quantum algorithms, which we would consider certain.
Note that a teleological character of the evolutions of the living on the Earth could be further explored by investigating, either theoretically or by means of toy models, the character of the operation of “tracing out all the rest”. We should consider the unitary quantum evolution of a system endowed with quantum speedup, standing for the evolution of the universe toward life, and derive from it the almost autonomous evolution of a subsystem of it, standing for the evolution of life on Earth, by tracing out all the rest. We should look for a way of obtaining an evolution of the subsystem that is teleological (thus quantum) in the longer term and (about) classical in the shorter term. If this were possible, it would go in favor of the quantum cosmological reading of the Anthropic Principle we are dealing with. By the way, this could also be another perspective under which to address the problem of unifying quantum and classical theory.
Another question is how to describe a macroscopic evolution of life endowed with quantum speedup. Of course, classical physics would turn out to be unable to describe not only the processes of the microscopic world but also the teleological evolutions of the macroscopic world.
@@
Let us end the present section by recalling what, to our knowledge, are the key steps that led to the discovery of the quantum computational speedup. According to the present theory, this discovery would be even more fundamental than currently thought. The first step was showing that computation is possible in the quantum world and introducing the fundamental notion of quantum bit, by Finkelstein [21] in 1969; that was also the first time that quantum physics moved to the world of the complex, indeed to a cellular automaton model of the universe—this was in the context of his well known attempt to unify relativity and quantum theory. Then, there was the introduction of the notion of logically reversible computation by Bennett [22] in 1973, of course, a prerequisite for the building up of computations in unitary quantum evolutions. The discovery of the superior computational efficiency of the quantum world with respect to that of the classical world in 1982 is due to Feynman [23]. Still in 1982, there has been the introduction of the notion of conservative logic by Fredkin and Toffoli [24]: with it, we have the logically reversible gates needed to perform quantum computations. The seminal quantum algorithm that delivers a quantum speedup was conceived by Deutsch [25] in 1985.
@@ 4 Conclusions
The flip side of the destabilizing character of the work is its innovative character. Let us distinguish between what should be certain and what should be plausible. What should be certain: (i) The work provides a unified quantitative explanation of the quantum speedups that is still missing from mainstream quantum mechanics. (ii) It provides a theoretical confirmation of the incompleteness of the quantum description of quantum nonlocality sustained by Einstein and others, also providing the local hidden variables envisaged by those authors to complete it. (iii) It highlights the teleological character of quantum algorithms; given the negative view that the teleological notion has in today’s science and the scope of this notion, this should be the most important result. What should be plausible: It shows that, under a quantum cosmological reading of the Fine-tuned Universe version of the Anthropic Principle, the physical basis of the teleological character of quantum algorithms applies as well to the evolutions of the living.
In conclusion, the present work should show that there is another possible view of physical reality out there, outside the current view. We note that this might not only be a curiosity. A revamping of the abandoned notion of the relation between evolution and functional structure of the living might have a profound impact on today’s evolutionary biology and biology.